OpenStudy (anonymous):

a reason lab is planning to explore the north pole. ten members of the lab each have a dog team. there are 56 dogs in all. if there are only 7- and 5- dog teams how many of each will there be ??

OpenStudy (anonymous):

By guess-and-check, three 7-dog teams and seven 5-dog teams.

OpenStudy (anonymous):

x = 5 dogs y = 7 dogs x + y = 10 5x + 7y = 56 y = 10 - x 5x + 70 - 7x = 56 -2x = -14 x = 7 y = 3 7 5-dog teams 3 7-dog teams